# FRAbench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v1

@@ D.2 Phased Training Experiment Results
Table 48: Average accuracy for phased training results on FRA-ID and FRA-ID-H test sets.
@@ D.3 GenEval-72B Experiment Results
We also train GenEval-72B using the same fine-tuning method as GenEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 15 and 16 and Tables 49 to 51, indicate that GenEval-72B offers some improvement over GenEval-7B. However, due to performance and deployment trade-offs, we find GenEval-7B to be more cost-effective.
Figure 15: Comparison GenEval-72B on FRA-ID and FRA-ID-H .
@@
Figure 16: Comparison GenEval-72B on FRA-OOD and FRA-OOD-H .
@@
Table 50: The results of GenEval-72B compared with baselines on three NLG benchmarks.
@@ Appendix E Experimental Details for Human Evaluation
We also compute the Inter-Annotator Agreement (IAA) for each annotator, with detailed results presented in Table 52 and Table 53. The final IAA among the three annotators achieved average scores of 84.2 and 87.4 for FRA-ID-H, and 84.0, 81.2, 78.2 for FRA-ODD-H, respectively. These results demonstrate a high level of reliability in our test dataset.
Annotators NLG IU IG ITIG
human12121-21 - 2 79.0 85.6 76.8 95.7
human13131-31 - 3 87.2 90.2 80.8 91.4
human23232-32 - 3 91.6 95.3 92.7 95.6
@@
Annotators NLG IU IG ITIG
human12121-21 - 2 80.3 95.0 76.1 84.7
human13131-31 - 3 88.1 81.3 73.9 82.4
human23232-32 - 3 77.7 79.1 77.4 78.6
@@ Appendix F The Prompt Template Used by UAs and TAs
In this section, we provide the prompt template used for training/inferencing GenEval. Note that We employed distinct prompt templates for UAs and TAs to enable specific aspect evaluation while avoiding overall assessment.
@@
Table 54: The example prompt templates for Universal Aspects only include the text output modality.
@@
Table 55: The example prompt templates for Universal Aspects only include the image output modality.
@@
Table 56: The example prompt templates for Universal Aspects only include the text-and-image output modality.
@@
Table 57: The example prompt templates for Task-specific Aspects in NLG task.
@@
Table 58: The example prompt templates for Task-specific Aspects in IU task.
@@
Table 59: The example prompt templates for Task-specific Aspects in IG task.
@@
Table 60: The example prompt templates for Task-specific Aspects in ITIG task with input.
@@
Table 61: The example prompt templates for Task-specific Aspects in ITIG task without input.
@@ Appendix G More Qualitative Examples
Table 62, Table 63, Table 64, Tables 65 to 67 show some good qualitative examples of feedback generated by GenEval. Table 68, Table 69, and Table 70 shows a relatively inferior example of feedback generated by GenEval.
@@
In the following example of Harmfulness Judgment, we speculate that the misunderstanding of GenEval is caused by over-associating some specific image feature with the criterion, yet it might be usable in some circumstances.
@@
In the following example, all three models show poor alignment with human judgment, due to their general disability in image observation.
@@
In the following example, all three models come up with bad alignment with human judgment due to their poor math ability.

### v2

@@ D.2 Staged Training Experiment Results
Table 51: Average accuracy for staged training results on FRA-ID and FRA-ID-H. Unified training obtains balanced or better performance.
@@ D.3 GenEval-72B Experiment Results
We also train GenEval-72B using the same fine-tuning method as GenEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 15 and 16 and Table 52 to 54, indicate that GenEval-72B offers some improvement over GenEval-7B. However, due to performance and deployment trade-offs, we find GenEval-7B to be more cost-effective.
Figure 15: Comparison GenEval-72B on FRA-ID and FRA-ID-H. We computed the average accuracy for both UAs and TAs across GPT-4o Alignment and Human Alignment conditions.
@@
Figure 16: Comparison GenEval-72B on FRA-OOD and FRA-OOD-H. “ow/unseen” denotes the average accuracy calculated only for unseen TAs. The aspect marked in blue is unseen TAs.
@@
Table 53: The results of GenEval-72B compared with different models on three NLG benchmarks.
@@ Appendix E Experimental Details for Human Evaluation
We also compute the Inter-Annotator Agreement (IAA) for each annotator, with detailed results presented in Table 55 and Table 56. The final IAA among the three annotators achieved average scores of 84.2, 87.4, and 93.8 for FRA-ID-H, and 84.0, 81.2, 78.2 for FRA-ODD-H, respectively. These results demonstrate a high level of reliability in our test dataset.
Annotator NLG IU IG ITIG
human $1-2$ 79.0 85.6 76.8 95.7
human $1-3$ 87.2 90.2 80.8 91.4
human $2-3$ 91.6 95.3 92.7 95.6
@@
Annotator NLG IU IG ITIG
human $1-2$ 80.3 95.0 76.1 84.7
human $1-3$ 88.1 81.3 73.9 82.4
human $2-3$ 77.7 79.1 77.4 78.6
@@ Appendix F The Prompt Template Used by UAs and TAs
In this section, we provide the prompt template used for training/inferencing GenEval in Table 57 to Table 64. Note that We employed distinct prompt templates for UAs and TAs to enable specific aspect evaluation while avoiding overall assessment.
@@
Table 57: The example prompt templates for UAs only include the text output modality.
@@
Table 58: The example prompt templates for UAs only include the image output modality.
@@
Table 59: The example prompt templates for UAs only include the multi-image output.
@@
Table 60: The example prompt templates for TAs in NLG task.
@@
Table 61: The example prompt templates for TAs in IU task.
@@
Table 62: The example prompt templates for TAs in IG task.
@@
Table 63: The example prompt templates for TAs in ITIG task with input.
@@
Table 64: The example prompt templates for TAs in ITIG task without input.
@@ Appendix G More Qualitative Examples
Table 65, Table 66, Table 67, Table 68 to 70 show four good qualitative examples of feedback generated by GenEval, corresponding to the NLG, IU, IG, and ITIG tasks, respectively. Table 71, Table 72, and Table 73 shows a relatively inferior example of feedback generated by GenEval.
@@
1. In the following example of Harmfulness Judgment, we speculate that the misunderstanding of GenEval is caused by over-associating some specific image feature with the criterion, yet it might be usable in some circumstances.
@@
2. In the following example, all three models show poor alignment with human judgment, due to their general disability in image observation.
@@
3. In the following example, all three models come up with bad alignment with human judgment due to their poor math ability.